OpenStudy (anonymous):

Bob needs to drive 592 miles to get to a family reunion. He travels 187 miles the first day. If he averages 52 miles per hour the second day, how long will it take him to reach his destination? Without calculating the answer, from the choices given, which is most reasonable? A. 20 hours B. 8 hours C. 3.75 hours D. 8 minutes

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

Total distance to be travelled = 592 miles DIstance already travelled in first day = 187 miles So, distance left to be travelled in the second day = 592 - 187 = 405 miles Average speed = 52 mph Time = Distance/Speed = 405/52 = 7.788462 hours

OpenStudy (anonymous):

@whpalmer4

OpenStudy (johnweldon1993):

I think @heather040200 had a good hold of this one :) Good work!

OpenStudy (anonymous):

Lol. Hoeeekaayyyy

OpenStudy (whpalmer4):

estimating, 597 miles - 187, call it 600-200 = 400. average speed of 52, call it 50. 400 / 50 is 8.

OpenStudy (anonymous):

sooo is a b c or d ???

OpenStudy (whpalmer4):

@valentina81 did you read any of the responses? Particularly mine and Heather's?